Evaluate the logarithm loga63 by using properties of logarithms and the following: loga3=5 loga7=2
Answer by venugopalramana(3286)

Evaluate the logarithm loga63 by using properties of logarithms and the following: loga3=5 loga7=2
SINCE ALL LOGS ARE TO BASE A LET ME DROP IT AND TAKE IT AS UNDERSTOOD,
FOR CONVENIENCE IN TYPING..
LOG(63)=LOG(7*3*3)=LOG(7)+LOG(3)+LOG(3)
=2+5+5=12